# Explain with Visual Key-Points Like a Real Mentor ! A Benchmark for Visual Solution-Explaining

### v1

@@ Abstract
With the rapid advancement of mathematical reasoning capabilities in large language models (LLMs), AI systems are increasingly being adopted in educational settings to support students’ comprehension of problem-solving processes. However, a critical component remains underexplored in current LLM-generated explanations: visual explanation. In real-world instructional contexts, human tutors routinely employ visual aids—such as diagrams, markings, and highlights—to enhance conceptual clarity. To bridge this gap, we introduce a novel task of visual solution explanation, which requires not only solving problems but also generating explanations that incorporate newly introduced visual elements essential for understanding (e.g., auxiliary lines, annotations, or geometric constructions). To evaluate model performance on this task, we propose MathExplain, a multimodal benchmark consisting of 997 math problems annotated with visual keypoints and corresponding explanatory text that references those elements. Our empirical results show that while some closed-source models demonstrate promising capabilities on visual solution-explaining, current open-source general-purpose models perform inconsistently, particularly in identifying relevant visual components and producing coherent keypoint-based explanations. We expect that visual solution-explaining and the MathExplain dataset will catalyze further research on multimodal LLMs in education and advance their deployment as effective, explanation-oriented AI tutors. Code and data will be released publicly.
@@ 1 Introduction
To address these limitations, we introduce visual solution-explaining, a novel task designed to evaluate and advance models’ ability to generate educationally effective, visually grounded mathematical explanations. To benchmark performance on visual solution-explaining, we propose multimodal Mathematics Explain (MathExplain) benchmark. The visual solution-explaining task requires identifying visual keypoints—elements not present in the original problem but crucial for understanding (e.g., lines, angles, annotations)—and generating explanatory text that explicitly refers to them. The MathExplain includes not only the original problem images and text but also solution images, explanatory text, and annotations of the visual keypoints that illustrate how the solution images differ from the originals (see Table 1). By providing textual representations of these visual keypoints, the benchmark allows even models—which currently struggle to generate fine-grained figures (Wu et al., 2024; Jo et al., 2024)—to capture and convey visually salient details in textual form. Notably, we emphasize the educational value of MathExplain, as it is derived from real-world instructional contexts.
@@ 2 Related Works
Recent advances in Large Language Models (LLMs) (Brown et al., 2020) have sparked significant interest in educational applications, particularly in personalized learning and automated tutoring (Park et al., 2024; Chevalier et al., 2024). These systems can adapt to individual learners by providing tailored feedback and customized curriculum (Hu et al., 2024; Macina et al., 2023; Feng et al., 2023). Beyond textual analysis, educational research indicates that providing both textual and visual information can more effectively bolster comprehension and memory than text alone (Arcavi, 2003; Stylianou, 2010; Lee et al., 2025). As Clark & Paivio (1991) notes, this advantage stems from the use of dual representations, which offer multiple retrieval cues and foster richer mental models for learners. To further harness these multimodal elements and provide more comprehensive learning opportunities, we introduce the visual solution-explaining task for an MLLM equipped with advanced multimodal understanding capabilities. This task enables the model to autonomously identify crucial visual keypoints—essential for understanding visual information—and integrate them into its explanations. By doing so, model can deliver more comprehensive educational support, ultimately enhancing the overall quality of student learning experiences.
@@
Current MLLMs excel at solving mathematical problems, making them highly useful tools for students (Zhuang et al., 2024; Luo et al., 2025). To evaluate these models, mathematical benchmarks (Cobbe et al., 2021; Hendrycks et al., 2021) have played a key role in assessing the reasoning capabilities of MLLMs. With the rise of multimodal models, benchmarks such as MathVista (Lu et al., 2023), Math-Vision (Wang et al., 2024), and MathVerse (Zhang et al., 2024) have extended these evaluations by introducing image-based math problems. Additionally, benchmarks like OlympiadBench (He et al., 2024) and MM-MATH (Sun et al., 2024) evaluate not only final answers but also the detailed solution process, providing deeper insights into model reasoning strategies. However, most existing benchmarks do not prioritize educational objectives. They generally focus on problem-solving performance rather than on the explanatory or pedagogical quality of solutions. Furthermore, the solutions they produce are predominantly textual, lacking the visual aids that can greatly enhance understanding. To address these gaps, we propose MathExplain, a new benchmark designed to assess a model’s visual solution-explaining capability—its ability to provide rich, multimodal explanations. By providing diverse and informative visuals, our benchmark advances educational research and supports the development of better teaching models.
@@ 3 MathExplain Benchmark
Figure 3: We propose two subtasks to robustly analyze visual solution-explaining capacity: (1) Visual Keypoint Identification, which challenges machines to generate visual keypoints useful for future explanation, and (2) Keypoint-based Explanation Generation, which requires models to generate explanations that explicitly reference the identified visual keypoints.
@@ 3.1 Benchmark Construction
We extract 997 instances of multimodal problem–solution pairs $\langle T_{p},\,I_{p},\,T_{s},\,I_{s}\rangle$ from an in-house mathematics education platform. All instances were authored by domain experts in mathematics to support effective student learning. The problems cover topics ranging from middle to high school level, primarily focusing on geometry and graph interpretation. All benchmark data were carefully curated to ensure compliance with copyright regulations.
To benchmark the models’ ability to recognize visual keypoints, we ensure that each solution image $I_{s}$ is derived from the corresponding problem image $I_{p}$ by adding only new elements—such as points, angles, lines, regions, and symbols—while preserving the original structure. For instance, in Figure 2, symbols and dashed lines are added to problem image $I_{p}$ to produce solution image $I_{s}$. This setup allows us to accurately evaluate whether a model can identify the critical visual keypoints and effectively incorporate them into its explanation.
For robust evaluation, we restrict the dataset to single-image math problems with either multiple-choice or short-answer formats. We focus on the domains of geometry and mathematical functions to ensure that visual context is essential for solving each problem. Each sample in MathExplain consists of two natural language texts (the problem text $T_{p}$ and the solution text $T_{s}$) and two RGB images (the problem image $I_{p}$ and the solution image $I_{s}$).
@@
To generate the visual keypoints (VK) in textual form, we use GPT-4o (Achiam et al., 2023) as an annotation tool. Identifying and formatting visual keypoints is a labor-intensive process; this tool assists human annotators by generating initial data frames. In our annotation workflow, the initial visual keypoints (VK) are generated by the AI tool and subsequently verified and refined by human annotators.
@@
    If any elements (points, lines, angles, regions, or symbols) are added or modified, they must be written in a {key: value} format. In this format, the key specifies the newly added or changed element, and the value describes how it is introduced by referencing surrounding elements.
@@
Following these principles, a set of the visual keypoints $\{vk^{AI}_{1},\dots,vk^{AI}_{n}\}\in VK^{AI}$ are generated by AI tools. After the model-generated results are produced, four human annotators with undergraduate or master’s degrees in science or engineering assess the quality of $VK^{AI}$ and check whether it adheres to the formatting guidelines. They then refine the $VK^{AI}$ to ensure both high quality and consistency with these principles. Lastly, the AI annotator translates the benchmark from Korean to English, and two human annotators verify and refine the translated content, ultimately forming the $VK$. The final structure of MathExplain is represented as $\langle T_{p},I_{p},T_{s},I_{s},VK\rangle$.
@@ 3.2 Data Analysis
The MathExplain benchmark consists of 997 problem–solution pairs: 765 (76.7%) geometry problems and 235 (23.5%) graph problems. Among these, 587 (58.8%) are multiple-choice questions and 423 (42.4%) are short-answer questions. It spans 17 chapters (see Figure 4) and 33 sections (see Appendix A). On average, each sample contains about 3.8 visual keypoints $VK$, derived from annotations. These keypoints fall into five main categories: points, lines, regions, and symbols. The symbol category is further divided into parallel marks, equal-length marks, right-angle marks, and length-label marks. Additional statistical details about $VK$, and length statistics for the problem text $T_{p}$, the solution text $T_{s}$, and the visual keypoint components $vk_{n}$ are provided in Table 2.
@@ 4 Task Definition
We propose two tasks to evaluate a model’s visual solution-explaining capability. visual solution-explaining requires identifying visual key-points useful for future explanation and generating explanations that explicitly reference them. These tasks offer a chance to assess how well the model understands multimodal educational elements and produces informative explanations along with visual keypoints. The two tasks are illustrated in Figure 3.
@@
The first task evaluates the model’s ability to infer visual keypoints that are crucial for comprehension. Given a problem image $I_{p}$ and problem text $T_{p}$, the model must infer the appropriate the visual keypoints $VK$ necessary for effective understanding. The $VK$ is represented as a list of dictionaries, where each key corresponds to a newly introduced visual element and the value provides a description of that element based on surrounding contextual features.
@@
The second task evaluates whether the model can effectively generate explanatory text by appropriately using the visual keypoints it has already recognized. Given a problem consisting of an image $I_{p}$ and text $T_{p}$, along with the visual keypoints $VK$, the model is required to produce a solution explanation $T_{s}$ that references the relevant visual elements.
@@ 5 Experiments
We evaluate both generalist and specialist MLLMs for multimodal mathematical problem solving. The generalist models include widely-used open-source MLLMs—LLaVA-1.6 (Liu et al., 2024), Molmo (Deitke et al., 2024), and Qwen-2.5-VL (Bai et al., 2025)—as well as proprietary models such as GPT-4o (Achiam et al., 2023) and Gemini 2.0 Flash (Google DeepMind, 2024). Specialist models focused on mathematics include Math-PUMA (Zhuang et al., 2024) and URSA (Luo et al., 2025). Details of the experimental setup are provided in Appendix A.
@@ 5.1 Visual Keypoint Identification
Given a problem image $I_{p}$ and accompanying text $T_{p}$, the model is expected to generate a set of visual keypoints $\{vk^{\text{Model}}_{1},\dots,vk^{\text{Model}}_{j}\}\in VK^{\text{Model}}$ that are essential for explaining the problem’s solution.
@@
We evaluate the model’s output by comparing the generated set $VK^{\text{Model}}$ with the ground-truth set $VK^{\text{GT}}=\{vk^{\text{GT}}_{1},\dots,vk^{\text{GT}}_{i}\}$ using exact match (EM) at the individual keypoint level. We report set-level precision ($EM_{P}$), recall ($EM_{R}$), and F1-score ($EM_{F1}$) as follows:
$EM_{P}=\frac{|VK^{\text{Match}}|}{|VK^{\text{Model}}|},\quad EM_{R}=\frac{|VK^{\text{Match}}|}{|VK^{\text{GT}}|},\quad EM_{F1}=2\times\frac{EM_{P}\times EM_{R}}{EM_{P}+EM_{R}}$
To establish matching between the two sets, two steps are required: (1) parsing the natural language model output into a structured set of discrete keypoint descriptions, and (2) determining whether each predicted keypoint semantically matches a ground-truth point. For parsing, we employ GPT‑4o (Achiam et al., 2023) to convert the generated text into JSON format. Each keypoint $vk_{n}$ is represented with an "element" field indicating the added visual component and a "description" field explaining how it modifies or extends the original content. Semantic matching is also performed using GPT‑4o, leveraging its strong language understanding capabilities. This allows for flexible alignment between corresponding elements despite label variations (e.g., a ground-truth “dotted line AB” matching a model output labeled “AZ”), enabling robust evaluation while accounting for surface-level discrepancies.
Figure 5: Failure cases of Qwen-2.5-VL and Math-PUMA in the Visual Keypoint Identification task. Qwen-2.5-VL shows ability to extract keypoints, whereas Math-PUMA mainly focuses on problem-solving and fails to identify them.
@@
As shown in Table 3, none of the current Open-source General (OG) or Open-source Mathematical (OM) models effectively identify visual keypoints. In contrast, Closed-source General (CG) models—Gemini 2.0 Flash and GPT-4o—demonstrate significantly stronger performance across all metrics, although they still exhibit certain limitations. Notably, both CG and OG models have higher $EM_{P}$ than $EM_{R}$, implying that while they capture some intuitive keypoints, they struggle with more detailed or complex ones. Despite being designed for mathematical reasoning, OM models perform unexpectedly poorly, because they often fail to follow instructions properly (see Figure 5). These models, with their strong focus on problem-solving, are not suitable for generating insightful explanations for students. In contrast, OG models at least make an effort to follow instructions when identifying visual keypoints. Furthermore, CG models not only follow instruction reliably but also maintain an understanding on task objectives, unlike OM models that narrowly focus on the problem-solving aspect.
@@ 5.2 Keypoint-based Explanation Generation
We assess explanation quality using three criteria: (1) Correctness – whether the model’s reasoning is logically sound and leads to a valid solution; (2) Fidelity – whether the explanation aligns with the reasoning and intent of the reference, regardless of surface form; (3) Visual Referencing – whether the explanation refers to the same key visual components (e.g., points, lines, angles, shapes) as the reference. Each criterion is rated on a 5-point Likert scale. We report annotation results from both human evaluators (Zheng et al., 2023) and an LLM-based evaluator (GPT‑4o (Achiam et al., 2023)). Complementary results based on conventional text similarity metrics (Papineni et al., 2002; Banerjee & Lavie, 2005) are provided in Appendix A.
@@
OM Math-PUMA 1.775 1.614 1.588 1.036 1.036 1.030
URSA 1.150 1.142 1.157 - - -
CG Gemini 2.0 Flash 3.872 3.508 4.126 4.361 4.060 4.084
@@
Table 4 shows that stronger, general-purpose models consistently outperform smaller, math-specific baselines, continuing the performance trend observed in the previous task. According to GPT-Eval, the closed-source model Gemini 2.0 Flash clearly leads across all three dimensions, Correctness, Fidelity, and Referencing. Among the open-source models, Qwen2.5-VL performs notably well, ranking third overall in GPT-Eval. In contrast, the math-specialist models struggle to produce logically consistent explanations, as reflected in their lower scores across all three dimensions. This underperformance is particularly striking for URSA, which previously excelled in mathematical benchmarks (Luo et al., 2025). Our qualitative analysis confirms that the specialist are optimized for problem solving rather than solution explaining. Human-Eval reveals similar findings. Gemini 2.0 Flash again receives the highest ratings, while Qwen2.5-VL continues to moderately outperform Math-PUMA. Overall, both GPT-Eval and Human-Eval reveal consistent trends—Gemini 2.0 Flash demonstrates a strong capacity to generate effective keypoint-based explanations, while most open-source models, especially those in the OM category, lag significantly behind. This highlights the performance gap between top-performing closed-source models and their open-source counterparts in visual solution-explaining tasks.

### v2

# Explain with Visual Keypoints Like a Real Mentor! A Benchmark for Multimodal Solution Explanation

@@ Abstract
With the rapid advancement of mathematical reasoning capabilities in Large Language Models (LLMs), AI systems are increasingly being adopted in educational settings to support students’ comprehension of problem-solving processes. However, a critical component remains underexplored in current LLM-generated explanations: visual explanation. In real-world instructional contexts, human tutors routinely employ visual aids—such as diagrams, markings, and highlights—to enhance conceptual clarity. To bridge this gap, we introduce a novel task of visual solution explanation, which requires generating explanations that incorporate newly introduced visual elements essential for understanding (e.g., auxiliary lines, annotations, or geometric constructions). To evaluate model performance on this task, we propose MathExplain, a multimodal benchmark consisting of 997 math problems annotated with visual keypoints and corresponding explanatory text that references those elements. Our empirical results show that while some closed-source models demonstrate promising capabilities on visual solution-explaining, current open-source general-purpose models perform inconsistently, particularly in identifying relevant visual components and producing coherent keypoint-based explanations. We expect that visual solution-explaining and the MathExplain dataset will catalyze further research on multimodal LLMs in education and advance their deployment as effective, explanation-oriented AI tutors. Code and data will be released publicly.
@@ 1 Introduction
To address these limitations, we introduce visual solution-explaining, a novel task designed to evaluate and advance models’ ability to generate educationally effective, visually grounded mathematical explanations. To benchmark performance on visual solution-explaining, we propose multimodal Mathematics Explain (MathExplain) benchmark. The visual solution-explaining task requires identifying visual keypoints that are not present in the original problem but crucial for understanding (e.g., lines, angles, annotations) and generating explanatory text that explicitly refers to them. The MathExplain includes not only the original problem images and text but also solution images, explanatory text, and annotations of the visual keypoints that illustrate how the solution images differ from the originals (see Table 1). By providing textual representations of these visual keypoints, the benchmark allows even models which currently struggle to generate fine-grained figures (Wu et al., 2024; Jo et al., 2024) to capture and convey visually salient details in textual form. Notably, we emphasize the educational value of MathExplain, as it is derived from real-world instructional contexts.
@@ 2 Related Works
Recent advances in Large Language Models (LLMs) (Brown et al., 2020) have sparked significant interest in educational applications, particularly in personalized learning and automated tutoring (Park et al., 2024; Chevalier et al., 2024). These systems can adapt to individual learners by providing tailored feedback and customized curriculum (Hu et al., 2024; Macina et al., 2023; Feng et al., 2023). Beyond textual analysis, educational research indicates that providing both textual and visual information can more effectively bolster comprehension and memory than text alone (Arcavi, 2003; Stylianou, 2010; Lee et al., 2025). As Clark & Paivio (1991) notes, this advantage stems from the use of dual representations, which offer multiple retrieval cues and foster richer mental models for learners. To further harness these multimodal elements and provide more comprehensive learning opportunities, we introduce the visual solution-explaining task for an Multimodal Large Language Models (MLLMs) equipped with advanced multimodal understanding capabilities. This task enables the model to autonomously identify crucial visual keypoints that are essential for interpreting visual information and integrate them into its explanations. By doing so, the model can provide more comprehensive educational support, ultimately enhancing the overall quality of student learning experiences.
@@
Current MLLMs excel at solving mathematical problems, making them highly useful tools for students (Zhuang et al., 2024; Luo et al., 2025). To evaluate these models, mathematical benchmarks (Cobbe et al., 2021; Hendrycks et al., 2021) have played a key role in assessing the reasoning capabilities of MLLMs. With the rise of multimodal models, benchmarks such as MathVista (Lu et al., 2023), Math-Vision (Wang et al., 2024), and MathVerse (Zhang et al., 2024) have extended these evaluations by introducing image-based math problems. Additionally, benchmarks like OlympiadBench (He et al., 2024) and MM-MATH (Sun et al., 2024) evaluate not only final answers but also the detailed solution process, providing deeper insights into model reasoning strategies. However, most existing benchmarks do not prioritize educational objectives. They generally focus on problem-solving performance rather than on the explanatory or pedagogical quality of solutions. Furthermore, the solutions they produce are predominantly textual, lacking the visual aids that can greatly enhance understanding. To address these gaps, we propose MathExplain, a new benchmark designed to assess a model’s visual solution-explaining capability, focusing on its ability to provide rich, multimodal explanations. By providing diverse and informative visuals, our benchmark advances educational research and supports the development of better teaching models.
@@ 3 MathExplain Benchmark
Figure 3: We propose two subtasks to robustly analyze visual solution-explaining capacity: (1) Visual Keypoint Identification, which challenges machines to generate visual keypoints useful for subsequent explanation, and (2) Keypoint-based Explanation Generation, which requires models to generate explanations that explicitly reference the identified visual keypoints.
@@ 3.1 Benchmark Construction
We extract 997 instances of multimodal problem–solution pairs $\langle T_{p},\,I_{p},\,T_{s},\,I_{s}\rangle$ from an in-house mathematics education platform. All instances were authored by domain experts in mathematics to support effective student learning. The problems cover topics ranging from middle to high school level, primarily focusing on domains of geometry and graph. All benchmark data were carefully curated to ensure compliance with copyright regulations.
To benchmark the models’ ability to recognize visual keypoints, we ensure that each solution image $I_{s}$ is derived from the corresponding problem image $I_{p}$ by adding only new elements such as points, angles, lines, regions, and symbols while preserving the original structure. For instance, in Figure 2, symbols and dashed lines are added to problem image $I_{p}$ to produce solution image $I_{s}$. This setup allows us to accurately evaluate whether a model can identify the critical visual keypoints and effectively incorporate them into its explanation.
For robust evaluation, we restrict the dataset to single-image math problems with either multiple-choice or short-answer formats. We focus on the domains of geometry and graph to ensure that visual context is essential for solving each problem. Each sample in MathExplain consists of two natural language texts (the problem text $T_{p}$ and the solution text $T_{s}$) and two RGB images (the problem image $I_{p}$ and the solution image $I_{s}$).
@@
To generate the visual keypoints $VK$ in textual form, we use GPT-4o (Achiam et al., 2023) as an annotation tool. Identifying and formatting visual keypoints is a labor-intensive process. This tool assists human annotators by generating initial data frames. In our annotation workflow, the initial visual keypoints $VK$ are generated by the AI tool and subsequently verified and refined by human annotators.
@@
    If any elements (points, lines, angles, regions, or symbols) are added or modified, they must be written in a {element: description} format. In this format, the ‘element’ specifies the newly added or changed element, and the ‘description’ indicates how it is introduced by referencing surrounding elements.
@@
Following these principles, a set of the visual keypoints $\{vk^{AI}_{1},\dots,vk^{AI}_{n}\}\in VK^{AI}$ are generated by AI tools. After $VK^{AI}$ are produced, four human annotators with undergraduate or master’s degrees in science or engineering assess its quality and verify its adherence to the formatting guidelines디. They then refine the $VK^{AI}$ to ensure both high quality and consistency with these principles. Lastly, the AI annotator translates the benchmark from Korean to English, and two human annotators verify and refine the translated content, ultimately forming the $VK$. The final structure of MathExplain is represented as $\langle T_{p},I_{p},T_{s},I_{s},VK\rangle$.
@@ 3.2 Data Analysis
The MathExplain benchmark consists of 997 problem–solution pairs: 765 (76.7%) geometry problems and 232 (23.3%) graph problems. Among these, 607 (60.9%) are multiple-choice questions and 390 (39.1) are short-answer questions. It spans 17 chapters (see Figure 4) and 33 sections (see Appendix B). On average, each sample contains about 3.8 visual keypoints $VK$, derived from annotations. These keypoints fall into five main categories: points, lines, regions, and symbols. The symbol category is further divided into parallel marks, equal-length marks, right-angle marks, and length-label marks. Additional statistical details about visual keypoints $VK$, and length statistics for the problem text $T_{p}$, the solution text $T_{s}$, and the visual keypoint components $vk_{n}$ are provided in Table 2.
@@ 4 Task Definition
We propose two tasks to evaluate a model’s visual solution-explaining capability. visual solution-explaining requires identifying visual keypoints useful for subsequent explanation and generating explanations that explicitly reference them. These tasks offer a chance to assess how well the model understands multimodal educational elements and produces informative explanations along with visual keypoints. The two tasks are illustrated in Figure 3.
@@
The first task evaluates the model’s ability to identify visual keypoints that are crucial for comprehension. Given a problem image $I_{p}$, problem text $T_{p}$, and problem’s answer, the model must infer the appropriate visual keypoints $VK$ necessary for an effective understanding. The $VK$ is represented as a list of JSON ({element: description}), where each ‘element’ corresponds to a newly introduced visual element and the ‘description’ provides contextual information about the element, referencing its surrounding content.
@@
The second task evaluates whether the model can effectively generate explanatory text using the appropriate visual keypoints it has already recognized. Given a problem consisting of an image $I_{p}$, text $T_{p}$, and problem’s answer, along with the visual keypoints $VK$, the model is required to produce a solution explanation $T_{s}$ that refers to the relevant visual elements.
@@ 5 Experiments
We evaluate both generalist and specialist MLLMs for multimodal mathematical problem solving. The generalist models include widely-used open-source MLLMs such as LLaVA-1.6 (Liu et al., 2024), Molmo (Deitke et al., 2024), and Qwen-2.5-VL (Bai et al., 2025), as well as proprietary models such as GPT-4o (Achiam et al., 2023) and Gemini 2.0 Flash (Google DeepMind, 2024). Specialist models focused on mathematics include Math-PUMA (Zhuang et al., 2024) and URSA (Luo et al., 2025). Details of the experimental setup and prompts are provided in Appendix A and C, respectively.
@@ 5.1 Visual Keypoint Identification
Given a problem image $I_{p}$ and accompanying text $T_{p}$ the model is expected to generate a set of visual keypoints $\{vk^{\text{Model}}_{1},\dots,vk^{\text{Model}}_{j}\}\in VK^{\text{Model}}$ that are essential for explaining the problem’s solution.
@@
We evaluate the model’s output by comparing the generated set $VK^{\text{Model}}$ with the ground-truth set $VK^{\text{GT}}=\{vk^{\text{GT}}_{1},\dots,vk^{\text{GT}}_{i}\}$ using Element Matching (EM) at the individual keypoint level. Let $VK^{\text{Match}}$ be the set of semantically matched keypoints between $VK^{\text{Model}}$ and $VK^{\text{GT}}$. We then report set-level precision $(EM_{P})$, recall $(EM_{R})$, and F1-score $(EM_{F1})$ as follows:
$EM_{P}=\frac{|VK^{\text{Match}}|}{|VK^{\text{Model}}|},\quad EM_{R}=\frac{|VK^{\text{Match}}|}{|VK^{\text{GT}}|},\quad EM_{F1}=2\times\frac{EM_{P}\times EM_{R}}{EM_{P}+EM_{R}}$ (1)
To determine $VK^{\text{Match}}$, we take two steps: (1) parsing the natural language model output into a structured set of discrete keypoint descriptions, and (2) determining whether each predicted keypoint semantically matches a ground-truth point. For parsing, we employ GPT‑4o (Achiam et al., 2023) to convert the generated text into JSON ({element: description}) format. Each keypoint $vk_{n}$ is represented with an ‘element’ field indicating the added visual component and a ‘description’ field explaining how it modifies or extends the original content. Semantic matching is also performed using GPT‑4o, leveraging its strong language understanding capabilities. This allows for flexible alignment between the corresponding elements despite label variations (e.g., a ground-truth ‘dotted line AB’ that matches a model output labeled ‘AZ’), enabling robust evaluation while accounting for surface-level discrepancies.
Figure 5: Failure cases of Qwen-2.5-VL and Math-PUMA in the Visual Keypoint Identification task. Qwen-2.5-VL shows ability to extract keypoints, whereas Math-PUMA mainly focuses solely on problem-solving and fails to identify them.
@@
As shown in Table 3, none of the current Open-source General (OG) or Open-source Mathematical (OM) models effectively identify visual keypoints. In contrast, Closed-source General (CG) such as Gemini 2.0 Flash and GPT-4o demonstrate significantly stronger performance across all metrics, although they still exhibit certain limitations. Notably, both CG and OG models have higher $EM_{P}$ than $EM_{R}$, implying that while they capture some intuitive keypoints, they struggle with more detailed or complex ones. Despite being designed for mathematical reasoning, OM models perform unexpectedly poorly, because they often fail to follow instructions properly (see Figure 5). These models, with their strong focus on problem-solving, are not suitable for generating insightful explanations for students. In contrast, OG models at least make an effort to follow instructions when identifying visual keypoints. Furthermore, CG models not only follow instruction reliably but also maintain an understanding on task objectives, unlike OM models that narrowly focus on the problem-solving aspect.
@@ 5.2 Keypoint-based Explanation Generation
We evaluate the quality of the explanation using three criteria: (1) Correctness – whether the model’s reasoning is logically sound and leads to a valid solution; (2) Fidelity – whether the explanation aligns with the reasoning and intent of the reference, regardless of surface form; (3) Visual Referencing – whether the explanation refers to the same key visual components (e.g. points, lines, angles, shapes) as the reference. Each criterion is rated on a 5-point Likert scale. We report annotation results from both human evaluators (Zheng et al., 2023) (see Appendix A for details) and an LLM-based evaluator (GPT‑4o (Achiam et al., 2023)). Complementary results based on conventional text similarity metrics (Papineni et al., 2002; Lin, 2004; Banerjee & Lavie, 2005; Zhang et al., 2019) are provided in Appendix A.
@@
OM Math-PUMA 1.499 1.164 1.061 1.041 1.041 1.037
URSA 1.107 1.079 1.132 - - -
CG Gemini 2.0 Flash 3.814 3.471 3.946 4.423 4.171 4.256
@@
GPT-Human ($\rho$)
Correctness 0.7695
Fidelity 0.7831
Referencing 0.7829
Table 5: Correlation ($\rho$) between the GPT and Human ratings.
@@
Table 4 shows that stronger, general-purpose models consistently outperform smaller, math-specific baselines, continuing the performance trend observed in the previous task. According to GPT-Eval, the closed-source model Gemini 2.0 Flash clearly leads across all three dimensions, Correctness, Fidelity, and Referencing. Among open-source models, Qwen2.5-VL performs notably well, ranking third overall in GPT-Eval. In contrast, math-specialist models struggle to produce logically consistent explanations, as reflected in their lower scores across all three dimensions. This underperformance is particularly striking for URSA, which previously excelled in mathematical benchmarks (Luo et al., 2025). We confirm that the specialists are optimized for problem-solving rather than solution-explaining. A similar trend is currently emerging in the Human-Eval results as well. As shown in Table 5, the Spearman correlation coefficient (Zar, 2005) is proving to be statistically significant. Gemini 2.0 Flash is once again receiving the highest ratings, while Qwen2.5-VL is continuing to moderately outperform Math-PUMA. Overall, both GPT-Eval and Human-Eval are revealing consistent trends: Gemini 2.0 Flash is demonstrating a strong capacity to generate effective keypoint-based explanations, while most open-source models, especially those in the OM category, are lagging significantly behind. This highlights the performance gap between top-performing closed-source models and their open-source counterparts in visual solution-explaining.
5.3 Does explanation quality reflect problem-solving ability?
For this analysis, ground-truth answers are excluded during inference to test if models can generate explanations without access to the correct answer. Models are first asked to generate an explanation without access to the correct answer (no-answer), and then asked to produce a hypothetical answer based solely on their self-generated explanation.
Result
Table 6 (left) presents GPT-Eval for the Keypoint-based Explanation Generation task, where models are asked to generate explanations without access to the final answer. Despite this constraint, the performance patterns remain broadly consistent with those observed in previous experiments. However, the absolute scores drop across all models, indicating that the complexity of explanation generation hinges on the availability of the correct answer.